National Accounts and Price Indices Consider an economy that only produces three types of goods and services: bread tractors, and haircuts. The following table shows how many of these goods were produced in three consecutive years and at what price they were sold. year bread tractors haircuts units sold price units sold price units sold price 2008 2009 2010 4,950,000 5,000,000 5,050,000 $1.00 S1.00 S1.05 95 100 95 $45,000 $50,000 $45,000 96,000 100,000 104,000 S50 $55 a) Calculate nominal GDP for these three years, as well as its growth rates b)Find real GDP at 2009 prices for the three years, as well as its growth rates c) Calculate the GDP deflator for the three years, as well as the inflation rates d) How are real GDP growth, nominal GDP growth, and the growth rate of the GDP e) Find the growth rate of real GDP using a chain-weighted GDP measure (see According to the statistics agency of this country, a typical household consumes 90 units implied by the GDP deflator. deflator related? Explain! textbook!) of bread and 2 haircuts per year f) Calculate the CPI implied by this goods bundle for the three years. What are the CPI inflation rates? g)Which of the growth rates you have calculated (or what combination thereof) best describes i. the change of economic activity, 11. the loss of value of the currency as felt by typical consumers, iii. the change of the number of consumption goods a typical household carn afford to buy, and iv. the change of the total dollar value of all transactions that take place in the economy? Briefly explain your answers (max. one sentence each)
Answer #1

(a) Nominal GDP (NGDP) = \sum(Current year price x Current year quantity)

2008 = 1 x 4,950,000 + 45,000 x 95 + 50 x 96,000 = 4,950,000 + 4,275,000 + 4,800,000 = 14,025,000

2009 = 1 x 5,000,000 + 50,000 x 100 + 50 x 100,000 = 5,000,000 + 5,000,000 + 5,000,000 = 15,000,000

2010 = 1.05 x 5,050,000 + 45,000 x 95 + 55 x 104,000 = 5,302,500 + 4,275,000 + 5,720,000 = 15,297,500

% change, 2008-2009 = (15,000,000 / 14,025,000) - 1 = 1.0695 - 1 = 0.0695 = 6.95%

% change, 2009-2010 = (15,297,500 / 15,000,000) - 1 = 1.0198 - 1 = 0.0198 = 1.98%

(b) Real GDP (RGDP) = \sum(Base year (2009) price x Current year quantity)

2008 = 1 x 4,950,000 + 50,000 x 95 + 50 x 96,000 = 4,950,000 + 4,750,000 + 4,800,000 = 14,500,000

2009 = 1 x 5,000,000 + 50,000 x 100 + 50 x 100,000 = 5,000,000 + 5,000,000 + 5,000,000 = 15,000,000

2010 = 1 x 5,050,000 + 50,000 x 95 + 50 x 104,000 = 5,050,000 + 4,750,000 + 5,200,000 = 15,000,000

% change, 2008-2009 = (15,000,000 / 14,500,000) - 1 = 1.0345 - 1 = 0.0345 = 3.45%

% change, 2009-2010 = (15,000,000 / 15,000,000) - 1 = 1 - 1 = 0%

(c) GDP Deflator = (NGDP / RGDP) x 100 and Inflation rate = % Change in GDP Deflator

2008 = (14,025,000 / 14,500,00) x 100 = 96.72

2009 = (15,000,000 / 15,000,000) x 100 = 100

2010 = (15,297,500 / 15,000,000) x 100 = 101.98

Inflation, 2008-2009 = (100 / 96.72) - 1 = 1.0339 - 1 = 0.0339 = 3.39%

Inflation, 2009-2010 = (101.98 / 100) - 1 = 1.0198 - 1 = 0.0198 = 1.98%

(d) Since GDP Deflator = (NGDP / RGDP) x 100,

% Change in GDP Deflator = % Change in NGDP - % Change in RGDP

Therefore, the higher (lower) the % change in NGDP and/or the lower (higher) the % change in RGDP, the higher (lower) the % change in GDP deflator.

NOTE: As per Answering Policy, 1st 4 parts are answered.